OpenStudy (anonymous):

Series expansion of 1/(z^2 + z + 1) about point 1+i

OpenStudy (anonymous):

\(\frac{1}{z^{2}+z+1}\) about \(z_{0}\) = \(1+i\)

OpenStudy (anonymous):

I'm assuming there's a way to do this without painstakingly using the definition, but I'm not quite sure how. I'm perfectly fine getting the expansion about \(z_{0} = 0\), just not sure what to do different in order to center it about 1 + i

ganeshie8 (ganeshie8):

is it possible to factor \(z^2+z+1\) and mimic the work for generating function of fibonacci series ?

ganeshie8 (ganeshie8):

perhaps this might help \[z^2+z+1 = \dfrac{z^3-1}{z-1}\]

ganeshie8 (ganeshie8):

\[\dfrac{1}{z^2+z+1}= \dfrac{1}{1-z^3} - \dfrac{z}{1-z^3}\]

OpenStudy (anonymous):

That definitely makes it look a lot easier (still would have to try it). How does this mimic the work for generating the fibonacci series? I havent really messed with anythin fibonacci

ganeshie8 (ganeshie8):

*fibonacci sequence

ganeshie8 (ganeshie8):

you may skip to 5th minute

OpenStudy (anonymous):

Alright, ill watch that first then see what I can come up with.

ganeshie8 (ganeshie8):

\(\large e^{\pm i2\pi/3}\) are the roots of \(z^2+z+1\), call these roots \(a,b\) respectively. so we can factor \(z^2+z+1\) as \((z-a)(z-b)\)

ganeshie8 (ganeshie8):

\[\dfrac{1}{z^2+z+1}= \dfrac{1}{(z-a)(z-b)}=\dfrac{1}{b-a}\left[\dfrac{1}{a-z} - \dfrac{1}{b-z}\right]\]

ganeshie8 (ganeshie8):

ofcourse hoping to use the geometric series \[\dfrac{1}{1-z} = \sum\limits_{n=0}^{\infty}z^n\]

OpenStudy (anonymous):

So from there, just do the appropriate manipulation? For example: \[\frac{ 1 }{ a-z } = \frac{ 1 }{ a-(1+i)-[z-(1+i)] } =\frac{ 1 }{ a-(1+i) }\cdot \frac{ 1 }{ 1-\frac{ z-(1+i) }{ a-(1+i) } }\]

ganeshie8 (ganeshie8):

that looks neat! honestly i never did these before, im just lifting my knowledge about "real" to "complex"

OpenStudy (anonymous):

Yeah, these turn out pretty cool, just I didnt consider making the roots into an arbitrary a and b. If Id have done that itd have come out much easier than me trying to integrate to arctan(z) and then differentiate that series back, lol. But yeah, its interesting getting more into these taylor and laurent series expansions. Thanks for the help and the idea ^_^

ganeshie8 (ganeshie8):

Awesome!